OpenStudy (anonymous):

sin 0 may be correctly expressed as which of the following? (all of these zeros of stright lines through them) a. cot 0/csc 0 b. cos 0 sec 0 c. csc 0 cot 0 d. tan 0 / sec 0

OpenStudy (anonymous):

@aum

OpenStudy (aum):

sin(x) = sin(x) / cos(x) * cos(x) = tan(x) * cos(x) = tan(x) / sec(x)

OpenStudy (micahm):

The Pythagorean trigonometric identity is a trigonometric identity expressing ... Values of 0 and π/2 are trivially proven by direct evaluation of sin and cos at those angles. ... By the periodicity identities we can say if the formula is true for −π < θ ≤ π ... The following table shows how this is done by dividing each element

OpenStudy (anonymous):

Everything you posted showed up as question marks :o @micahm which one is it..

OpenStudy (micahm):

A or B is what i had in mind

OpenStudy (aum):

In my previous reply I showed sin(x) = tan(x) / sec(x). Therefore, sin(0) = tan(0) / sec(0)

OpenStudy (anonymous):

Oh I see